What is Joule-Thomson effect ? Obtain the expression for the Joule-Thomson coefficient for a real gas.
1
Expert's answer
2020-06-08T10:14:13-0400

Answer:

In thermodynamics, the Joule-Thomson effect (also known as the Joule-Kelvin effect, the Kelvin-Joule effect) describes the temperature change of a real gas or liquid (differentiated from an ideal gas) when it is passed through a valve or porous plug, while keeping it isolated. so that heat does not exchange with the environment. This procedure is called the throttling process or the Joule-Thomson process. At room temperature, all gases except hydrogen, helium and cool neon are expanded by the Joule-Thomson process when throttled through an opening; these three gases experience the same effect, but only at lower temperatures. Most fluids such as hydraulic oils will heat up during the Joule-Thomson throttling process.

The rate of change of temperature with respect to pressure in the Joule-Thomson process (that is, with constant enthalpy) is the Joule-Thomson (Kelvin) coefficient.

"\mu_{JT}=(\dfrac{dT}{dP})_H=\dfrac{V}{C_p}(\alpha T-1)."

The Joule-Thomson coefficient is measured in K/Pa.